Plsss make is correct! ! Written clearly!!


1-56 An airline in a small city has five departures each day. It is known that any given flight has a probability of 0.3 of departing late. For any given day find the probabilities that: (a) no flights depart late, (b) all flights depart late, and (c) three or more depart on time.
media%2F74c%2F74c5070c-49eb-4902-8983-8f
Answer #1

Question 1.56) XBinomial(n -5,p 0.3) PlX-0)-(2)pr(1 _ p)--(o) (0.3)(1-0.3)5-0-1 * (0.3)(0.7) 0.16807 0.168070 0.168 P(X 0C P(Y > 3), n-5, p-0.7 = P(Y = 3) + P(Y-4) + P(Y-5) (because these are disjoint events) PY 3 () (0.7)3(1 -0.7)2 10(0.7) (0.3)